OpenStudy (anonymous):

Factor the polynomial and find all (real and fake) its roots: x^5-4x^3-x^2+4

OpenStudy (anonymous):

f(1) = 1-4-1 + 4 = 0 so (x-1) is a factor we can now divide by x-1 and se if the quotient can be factored but because of the +4 at the end 2 or -2 may be factors: f(2) = 32 - 32 - 4 + 4 = 0 f(-2) = = -32 + 32 - 4 + 4 = 0 so (x+2) and (x -2) are also factors

OpenStudy (anonymous):

i have to go now - there are another 2 factors which could be complex (what u called fake?)

OpenStudy (anonymous):

sorry - must go

OpenStudy (anonymous):

Yes but go you know how you would divide those factors into the original equation. And yes, i call them fake, which they are imagery, or complex, which comes in pairs and are Complex Conjugates of each other.

OpenStudy (mertsj):

This is easier than I thought. Let's try factoring by grouping: x&5

OpenStudy (mertsj):

(x^5-4x^3)-(x^2-4) x^3(x^2-4)-1(x^2-4) (x^3-1)(x^2-4) (x-1)(x^2+x+1)(x-2)(x+2) x=1,2,-2 and the complex roots of x^2+x+1

OpenStudy (mertsj):

Use the quadratic formula to find the roots of x^2+x+1

OpenStudy (anonymous):

ok i can do that